2 cubed-root(a^2) / (square-root of 8a)... Im Having trouble with this problem. -thx for your help
Answer by fractalier(6550)

Okay we have
2*cbrt(a^2) / sqrt(8a) =
2*cbrt(a^2 / 2*sqrt(2a) =
cbrt(a^2) / sqrt(2)*sqrt(a) =
a^(2/3) / sqrt(2)*a^(1/2) = remember that 2/3 - 1/2 = 1/6
a^(1/6) / sqrt(2) now rationalize the denominator and get
sqrt(2)*a^(1/6) / 2